OpenStudy (anonymous):

simplify 3 ^3(sqrt)24?

OpenStudy (saifoo.khan):

\[\Large \sqrt[3]{24}\]

OpenStudy (saifoo.khan):

like this?

OpenStudy (anonymous):

Yes.

OpenStudy (saifoo.khan):

Factor 24.

OpenStudy (anonymous):

8 * 8 * 8 so 8^3

OpenStudy (saifoo.khan):

No. 8*3

OpenStudy (saifoo.khan):

Now factor further.

OpenStudy (anonymous):

2*2*2*3?

OpenStudy (anonymous):

??

OpenStudy (anonymous):

ok....

OpenStudy (saifoo.khan):

Yes.

OpenStudy (anonymous):

what do i do now?

OpenStudy (saifoo.khan):

cuberoot(24) = cuberoot(8*3) = cuberoot(2*2*2*3) = 2cuberoot8

OpenStudy (anonymous):

\[a. 6\sqrt[3]{3}\] b. 18 \[c. 9\sqrt[3]{3}\] d. 27